# LLM as a Broken Telephone: Iterative Generation Distorts Information

### v1

@@ Abstract
Amr Mohamed1, Mingmeng Geng2, Michalis Vazirgiannis1,3, Guokan Shang1 1MBZUAI, 2SISSA, 3Ecole Polytechnique {amr.mohamed, guokan.shang}@mbzuai.ac.ae
@@ 1 Introduction
This prompts us to consider the question: what effects arise when the same piece of information is repeatedly processed by LLMs through multiple iterations? This procedure is analogous to the telephone game in human communication, a widely known children’s game in which a message is passed sequentially from one player to the next, with the final version often differing significantly from the original, usually with amusing or humorous effect. This happens because players often act as broken telephones, where information is gradually distorted as it is passed along the chain of individuals, highlighting how repeated transmission can lead to the accumulation of errors, omissions, or unintended alterations (Hitchcock et al., 2011).
@@
In our study, we simulate the LLMs’ telephone game in conjunction with the translation task mainly, under three experimental setups. As illustrated in Figure 1, within each iteration, a document in English is subsequently translated into one or more different languages, then back to English, by leveraging LLMs. We compare the back-translated English version with the initial English version at every iteration with textual relevance and factuality measures, to investigate whether and how information distortion accumulates. Our results show that over time, small alterations in phrasing, meaning, or factual details can accumulate, leading to a progressive drift from the original source, as illustrated by the example in Table 1. Code and data are publicly available [^1] https://github.com/amr-mohamedd/LLM-as-a-Broken-Telephone. Our main findings include:
@@ 3.1 Notations and Definitions
Let $\mathcal{D}=\{d_{i}\}_{i=1}^{I}$ denote a set of $I$ documents, =$\mathcal{L}=\{l_{j}\}_{j=1}^{J}$ as a set of $J$ natural languages, and =$\mathcal{M}=\{m_{k}\}_{k=1}^{K}$ for a set of $K$ models.
We define a translation chain as a sequence of $N$ translation iterations that progressively transform a document. For iteration $t\geq 1$, let $d_{i,l_{\text{source}}}^{(t-1)}$ be the $i$-th document in the source language at iteration $t-1$. At iteration $t$, an ordered language chain is constructed by selecting a permutation $\pi^{(t)}$ of $J-1$ languages from $\mathcal{L}$ and forming the sequence
@@ 3.2 Experimental Settings
Decoding Parameters and Translation Prompt. Each model was used for inference with its default decoding parameters. We capped the maximum number of newly generated tokens at 8000 to encourage open-ended generation. This high limit allows translations, which can vary in length across different languages, to conclude naturally rather than being prematurely truncated. Models within the main experiments were prompted to translate documents from a source to a target language with a moderately constrained prompt. The full translation prompt can be found in Appendix C.
@@ 3.3 Evaluation Metrics
Factuality Preservation. FActScore (Min et al., 2023) decomposes long-form text into atomic units and verifies each against a trusted reference using a dedicated judge model. In this study, we assume that the original text is factually correct and use FActScore to assess the rate of factuality degradation over the different iterations by comparing each model generation with its original text, then employ Claude 3.5 Sonnet as the judge model.
@@ 4.1 Experiment 1: Bilingual Self-loop
$\mathcal{L}$ =$=\{\text{EN},\,l_{\text{bridge}}\}$ (5)
@@
$\;\rightarrow\;l_{\text{bridge}}\;\rightarrow\;\text{EN}.$ EN . (6)
@@
[…]
@@ 4.2 Experiment 2: Bilingual Two-player
$\mathcal{L}$ =$=\{\text{EN},\,l_{\text{bridge}}\}$ (9)
@@
[…]
@@ 4.3 Experiment 3: Multilingual Multiplayer
=$\mathcal{L}=\{\text{EN},\text{FR},\text{TH}\}\quad$ (12)
@@
$\rightarrow l_{1}^{(t)}\rightarrow l_{2}^{(t)}\rightarrow\text{EN},$ EN ,
[…]
@@
=$\mathcal{L}=\{\text{EN},\text{FR},\text{TH},\text{ZH},\text{DE}\},$ (16)
@@
Results. As shown in Appendix B.2, all three experimental settings indicated a comparable degree of factual, lexical, and semantic degradation by the $100^{th}$ iteration across all evaluation metrics. However, differences emerged in the rate at which this degradation occurred. Specifically, Setting 3 exhibited the steepest decline in factual accuracy, with an average FActScore gradient of $-0.038\pm 0.02$. By the $10^{th}$ generation, Setting 3’s factuality had dropped to $0.054$, and by the $100^{th}$ generation, it further declined to $0.04$. Setting 1 followed closely, with an average gradient of $-0.036\pm 0.02$, showing a factuality score of $0.063$ at the $10^{th}$ generation and $0.04$ at the $100^{th}$. Setting 2 exhibited the slowest rate of factual degradation, with an average gradient of $-0.034\pm 0.02$, reaching $0.075$ at the $10^{th}$ generation and $0.04$ at the $100^{th}$.
@@ 5 Ablation Studies
[…]
@@ 5.1 Other Tasks: Rephrasing
In this task, the model is instructed to rephrase a given document while preserving its full meaning (the full rephrasing prompt can be found in Appendix C). We randomly sampled 30 documents from News2024 and conducted four experiments based on the setups from Sections 4.1, 4.2, and 4.3 (Setting 2). These experiments tested standalone rephrasing chains, the collaborative effects of Llama and Mistral, and an extended setup incorporating Gemma into the chain.
@@ 5.2 Temperature Variation Affects Outputs
From Figure 5, higher temperature settings lead to greater factual and semantic degradation. At extremely low temperatures ($1\times 10^{-6}$), factuality drops slightly in the first two iterations but stabilizes thereafter. As temperature increases, stability diminishes, and factuality gradually diverges. Higher temperatures exacerbate this trend, with maximum temperature (1.0) causing the steepest decline, showing continuous divergence. Additional examples can be found in Appendix D.
@@ 5.3 Sensitivity of Iterative Translation Outputs to the Chosen Prompt
We subsequently investigated the influence of the translation prompt on the outputs produced by the iterative process. To this end, 30 documents were randomly sampled from the News2024 dataset, and Llama was tasked with translating them using three distinct prompts characterized by varying levels of constraint: simple, base (used in all our experiments), and constrained. The complete prompts are provided in Appendix C.
@@ 6 Discussion and Conclusion
While our study uses datasets from three different domains—book summaries, movie scripts, and news—they share similar characteristics and may not capture the rare or long-tailed information typical of specialized domains. Additionally, due to computational resource constraints, our experiments focus on models with seven to nine billion parameters. Future work should explore whether datasets from specialized domains and larger models influence the levels of information distortion in iterative generation chains.
@@ B.2 Experiment 3: Multilingual Multiplayer
Appendix C Generation Prompts
@@
Appendix D Examples Analysis

### v2

@@ Abstract
Amr Mohamed1, Mingmeng Geng2, Michalis Vazirgiannis1,3, Guokan Shang1 1MBZUAI, 2SISSA, 3Ecole Polytechnique Correspondence: {amr.mohamed, guokan.shang}@mbzuai.ac.ae
@@ 1 Introduction
This prompts us to consider the question: what effects arise when the same piece of information is repeatedly processed by LLMs through multiple iterations? This procedure is analogous to the telephone game in human communication, a widely known children’s game in which a message is passed sequentially from one player to the next, with the final version often differing significantly from the original, usually with amusing or humorous effect. This happens because players often act as broken telephones, where information is gradually distorted as it is passed along the chain of individuals, highlighting how repeated transmission can lead to the accumulation of errors, omissions, or unintended alterations (Whallon et al., 2011).
@@
In our study, we simulate the LLMs’ telephone game primarily through the task of translation, as iterative translation serves as a critical and tangible testbed for examining how meaning and form degrade over repeated generations. This setup reflects real-world scenarios—for instance, cross-lingual news transmission—where content is repeatedly translated across languages. The extent of distortion, shaped by the interplay between a language’s representation in the training data and its linguistic similarity to the source, is not unique to translation but illustrates a broader phenomenon also observed in other iterative LLM tasks, such as rephrasing, which we investigate under three experimental setups.
As illustrated in Figure 1, within each iteration, a document in English is subsequently translated into one or more different languages, then back to English, by leveraging LLMs. We compare the back-translated English version with the initial English version at every iteration with textual relevance and factuality measures, to investigate whether and how information distortion accumulates. Our results show that over time, small alterations in phrasing, meaning, or factual details can accumulate, leading to a progressive drift from the original source, as illustrated by the example in Table 1. Code and data are publicly available [^1] https://github.com/amr-mohamedd/LLM-as-a-Broken-Telephone. Our main findings include:
@@ 3.1 Notations and Definitions
Let $\mathcal{D}=\{d_{i}\}_{i=1}^{I}$ denote a set of $I$ documents, $\mathcal{L}=\{l_{j}\}_{j=1}^{J}$ as a set of $J$ natural languages, and $\mathcal{M}=\{m_{k}\}_{k=1}^{K}$ for a set of $K$ models.
We define a translation chain as a sequence of $N$ translation iterations that progressively transform a document. For iteration $t\geq 1$, let $d_{i,l_{\text{source}}}^{(t-1)}$ be the $i$-th document in the source language at iteration t1t-1. At iteration $t$, an ordered language chain is constructed by selecting a permutation $\pi^{(t)}$ of J1J-1 languages from $\mathcal{L}$ and forming the sequence
@@ 3.2 Experimental Settings
Decoding Parameters and Translation Prompt. Each model was used for inference with its default decoding parameters. We capped the maximum number of newly generated tokens at 8000 to encourage open-ended generation. This high limit allows translations, which can vary in length across different languages, to conclude naturally rather than being prematurely truncated. Models within the main experiments were prompted to translate documents from a source to a target language with a moderately constrained prompt. The full translation prompt can be found in Appendix E.
@@ 3.3 Evaluation Metrics
Factuality Preservation. FActScore (Min et al., 2023) decomposes long-form text into atomic units and verifies each against a trusted reference using a dedicated judge model. In this study, we assume that the original text is factually correct and use FActScore to assess the rate of factuality degradation over the different iterations by comparing each model generation with its original text, then employ Claude 3.5 Sonnet to be the judge model.
@@ 4.1 Experiment 1: Bilingual Self-loop
$\mathcal{L}$ $=\{\text{EN},\,l_{\text{bridge}}\}$ (5)
@@
$\text{EN}\;\rightarrow\;l_{\text{bridge}}\;\rightarrow\;\text{EN}.$ (6)
@@
The average gradient values of FActScore in Table 2 quantify the rate of factuality loss across translation iterations. For language pairs composed solely of Latin script languages, gradients remain close to zero across all datasets and LLMs, indicating minimal degradations. For instance, in the News2024 dataset, the average gradients for $\textbf{EN}\leftrightarrow\textbf{FR}$ are -0.004 (±0.003) with Llama and -0.007 (±0.004) with Mistral, while for $\textbf{EN}\leftrightarrow\textbf{DE}$ they are -0.005 (±0.003) and -0.011 (±0.006), respectively. In contrast, chains involving non-Latin scripts—particularly Thai—exhibit significantly faster factuality loss. In the BookSum dataset, the $\textbf{EN}\leftrightarrow\textbf{TH}$ gradient is -0.026 (±0.014) with Llama and -0.040 (±0.025) with Mistral. This pattern is consistently observed across all evaluated language pairs, datasets, and models, with Thai demonstrating the highest rates of factual degradation.
@@ 4.2 Experiment 2: Bilingual Two-player
$\mathcal{L}$ $=\{\text{EN},\,l_{\text{bridge}}\}$ (9)
@@
Figure 3: Comparison of metrics for the Bilingual Two-Player Experiment on the News2024 dataset, illustrating the interaction effects between Llama and Mistral on translation chains for $\text{EN}\leftrightarrow\text{FR}$ and $\text{EN}\leftrightarrow\text{TH}$, contrasted with their individual performances in the Bilingual Self-loop Experiment.
@@ 4.3 Experiment 3: Multilingual Multiplayer
$\mathcal{L}=\{\text{EN},\text{FR},\text{TH}\}\quad$ (12)
@@
$\text{EN}\rightarrow l_{1}^{(t)}\rightarrow l_{2}^{(t)}\rightarrow\text{EN},$
with $l_{1}^{(t)}$ and $l_{2}^{(t)}$ drawn from $\{\text{FR},\text{TH}\}$ and satisfying $l_{1}^{(t)}\neq l_{2}^{(t)}$. The corresponding model sequence is
@@
$\mathcal{L}=\{\text{EN},\text{FR},\text{TH},\text{ZH},\text{DE}\},$ (16)
@@
Results. As shown in Appendix B.2, all three experimental settings indicated a comparable degree of factual, lexical, and semantic degradation by the $100^{th}$ iteration across all evaluation metrics. However, differences emerged in the rate at which this degradation occurred. Specifically, Setting 3 exhibited the steepest decline in factual accuracy, with an average FActScore gradient of $0.038\pm 0.02$. By the $10^{th}$ generation, Setting 3’s factuality had dropped to $0.054$, and by the $100^{th}$ generation, it further declined to $0.04$. Setting 1 followed closely, with an average gradient of $0.036\pm 0.02$, showing a factuality score of $0.063$ at the $10^{th}$ generation and $0.04$ at the $100^{th}$. Setting 2 exhibited the slowest rate of factual degradation, with an average gradient of $0.034\pm 0.02$, reaching $0.075$ at the $10^{th}$ generation and $0.04$ at the $100^{th}$.
@@ 5 Ablation Studies
Figure 5: Impact of temperature variation on Llama outputs for 30 randomly sampled documents from the News2024 dataset, evaluated on the $\text{EN}\leftrightarrow\text{FR}$ translation chain.
@@ 5.1 Other Tasks: Rephrasing
In this task, the model is instructed to rephrase a given document while preserving its full meaning (the full rephrasing prompt can be found in Appendix E). We randomly sampled 30 documents from News2024 and conducted four experiments based on the setups from Sections 4.1, 4.2, and 4.3 (Setting 2). These experiments tested standalone rephrasing chains, the collaborative effects of Llama and Mistral, and an extended setup incorporating Gemma into the chain.
@@ 5.2 Temperature Variation Affects Outputs
From Figure 5, higher temperature settings lead to greater factual and semantic degradation. At extremely low temperatures ($1\times 10^{-6}$), factuality drops slightly in the first two iterations but stabilizes thereafter. As temperature increases, stability diminishes, and factuality gradually diverges. Higher temperatures exacerbate this trend, with maximum temperature (1.0) causing the steepest decline, showing continuous divergence. Additional examples can be found in Appendix F.
@@ 5.3 Sensitivity of Iterative Translation Outputs to the Chosen Prompt
We subsequently investigated the influence of the translation prompt on the outputs produced by the iterative process. To this end, 30 documents were randomly sampled from the News2024 dataset, and Llama was tasked with translating them using three distinct prompts characterized by varying levels of constraint: simple, base (used in all our experiments), and constrained. The complete prompts are provided in Appendix E.
@@ 6 Discussion and Conclusion
While our study utilizes datasets from three distinct domains—book summaries, movie scripts, and news articles—these sources share similar characteristics and may not reflect the rare or long-tailed information found in specialized domains. Moreover, due to computational resource limitations, our experiments are restricted to models with 7–9 billion parameters, using their default generation settings. Future work should investigate whether incorporating datasets from specialized domains, employing larger models, or varying generation strategies (e.g., greedy decoding) impacts the degree of information distortion in iterative generation chains.
@@ B.2 Experiment 3: Multilingual Multiplayer
Appendix C Generation Settings and Decoding Parameters
Each model was used for inference with its default decoding parameters as specified in their respective official documentation. We capped the maximum number of newly generated tokens at 8000 to encourage open-ended generation and ensure translations could conclude naturally. The choice of default parameters was made to reflect common practical usage scenarios where models are often deployed with these settings. The default sampling-based decoding parameters used for each model in our experiments are detailed in Table 3.
Model Temperature Top-p
Llama-3.1-8B-Instruct 0.6 0.9
Mistral-7B-Instruct-v0.2 0.0 N/A
Gemma-2-9B-it 1.0 0.95
Table 3: Default decoding parameters used for LLama, Mistral, and Gemma.
Further exploration of diverse decoding strategies, such as greedy decoding or beam search, remains an avenue for future work.
Appendix D Additional BLEURT Score Evaluations
To further reinforce our findings, we computed BLEURT scores (Sellam et al., 2020)—known for their high correlation with human judgments—across the full outputs of our primary experiments and rephrasing tasks. While main evaluations assessed FActScore for factual preservation alongside traditional textual relevance metrics (BLEU, ROUGE, METEOR, BERTScore) to capture changes in linguistic form over successive iterations, the BLEURT scores presented below (Tables 4 to 7) further substantiate our conclusions regarding information distortion. Higher BLEURT scores indicate better quality and closer semantic similarity to the original text.
D.1 Bilingual Self-loop BLEURT Scores
Table 4 presents the BLEURT scores for the Bilingual Self-loop experiment, corresponding to the Llama model on the News2024 dataset (as detailed in Section 4.1 and visualized for other metrics in Figure 2).
Language Pair/ Iteration
Iteration 0 3 6 10 15 20 30 40 50 65 80 100
EN $\leftrightarrow$ FR 0.949 0.727 0.714 0.704 0.696 0.693 0.688 0.683 0.678 0.676 0.675 0.670
EN $\leftrightarrow$ NL 0.949 0.721 0.704 0.689 0.679 0.670 0.664 0.657 0.654 0.649 0.645 0.642
EN $\leftrightarrow$ DE 0.949 0.711 0.688 0.677 0.669 0.660 0.654 0.645 0.641 0.635 0.630 0.625
EN $\leftrightarrow$ VN 0.949 0.681 0.659 0.643 0.632 0.625 0.614 0.606 0.603 0.596 0.589 0.586
EN $\leftrightarrow$ ZH 0.949 0.618 0.587 0.570 0.551 0.537 0.522 0.514 0.508 0.498 0.494 0.489
EN $\leftrightarrow$ TH 0.950 0.584 0.537 0.514 0.496 0.482 0.462 0.454 0.447 0.440 0.434 0.426
Table 4: BLEURT scores for the Bilingual Self-loop experiment using Llama on the News2024 dataset. Scores show the evolution of text quality over 100 iterations for different language pairs.
Similar to our previous results for textual relevance and factuality in this experimental setup, the BLEURT scores exhibit the same trends. Specifically, there is a consistent decline in scores across all iterations, indicating progressive degradation of text quality. This degradation is less severe for language pairs where the bridge language uses a Latin script and shares more similarities with English (e.g., EN $\leftrightarrow$ FR, EN $\leftrightarrow$ NL, EN $\leftrightarrow$ DE), which show higher BLEURT scores throughout the iterations compared to pairs involving non-Latin scripts or more distant languages (e.g., EN $\leftrightarrow$ VN, EN $\leftrightarrow$ ZH, and particularly EN $\leftrightarrow$ TH). This observation aligns with Hypothesis 1, which posited that lexical and script similarity would influence the degree of information distortion.
D.2 Bilingual Two-Player BLEURT Scores
Table 5 shows BLEURT scores for the Bilingual Two-Player experiment on the News2024 dataset (detailed in Section 4.2 and Figure 3), involving Llama and Mistral.
Language Pair/ Iteration
Iteration 0 3 6 10 15 20 30 40 50 65 80 100
EN $\leftrightarrow$ FR 0.949 0.684 0.661 0.640 0.622 0.614 0.597 0.592 0.575 0.570 0.564 0.547
EN $\leftrightarrow$ TH 0.949 0.392 0.351 0.330 0.327 0.318 0.288 0.284 0.270 0.271 0.262 0.264
Table 5: BLEURT scores for the Bilingual Two-Player experiment on the News2024 dataset.
D.3 Multilingual Multiplayer BLEURT Scores
Table 6 shows BLEURT scores for the Multilingual Multiplayer experiment on the News2024 dataset (detailed in Section 4.3 and Appendix B.2, Figure 12).
Setting / Iteration
Iteration 0 3 6 10 15 20 30 40 50 65 80 100
Setting 1 0.949 0.385 0.348 0.334 0.333 0.329 0.311 0.295 0.305 0.304 0.302 0.295
Setting 2 0.949 0.414 0.360 0.330 0.328 0.312 0.312 0.292 0.284 0.275 0.274 0.283
Setting 3 0.949 0.397 0.350 0.326 0.312 0.307 0.309 0.303 0.298 0.296 0.287 0.300
Table 6: BLEURT scores for the Multilingual Multiplayer experiment on the News2024 dataset.
D.4 Rephrasing Task BLEURT Scores
Table 7 shows BLEURT scores for the rephrasing task on the News2024 dataset (detailed in Section 5.1 and Figure 4).
Model Combination / Iteration
Iteration 0 3 6 10 15 20 30 40 50 65 80 100
Llama 0.945 0.645 0.619 0.602 0.591 0.587 0.580 0.567 0.557 0.551 0.542 0.537
Mistral 0.945 0.593 0.577 0.562 0.552 0.545 0.538 0.535 0.533 0.533 0.528 0.521
Llama + Mistral 0.945 0.614 0.585 0.572 0.558 0.548 0.538 0.518 0.512 0.502 0.492 0.490
Llama + Mistral + Gemma 0.945 0.587 0.552 0.531 0.507 0.496 0.480 0.465 0.457 0.443 0.438 0.427
Table 7: BLEURT scores for the rephrasing task on the News2024 dataset.
These BLEURT evaluations, due to their high alignment with human judgment, reinforce our original analysis and confirm that the pattern of iterative degradation observed in our experiments is robust. The decline in BLEURT scores over iterations across different tasks and settings provides further evidence for the "broken telephone" effect in LLM-based iterative generation.
Appendix E Generation Prompts
@@
Appendix F Examples Analysis